OpenStudy (anonymous):

How do you verify the identity of the following : ((cosx)/(1+sinx))+((1+sinx)/(cosx))=2secx PLEASE HELP!

OpenStudy (anonymous):

@mayaal any chance you could help?

OpenStudy (mayaal):

yup

OpenStudy (mayaal):

1+ sinx)/(cos x) + (cos x)/(1+sin x) =(1+ sinx)/(cos x) + [(cos x)(1-sinx)]/[(1+sin x)(1-sinx)] =(1+ sinx)/(cos x) + [(cos x)(1-sinx)]/(1-sin²x) =(1+ sinx)/(cos x) + [(cos x)(1-sinx)]/(cos²x) =(1+ sinx)/(cos x) + (1-sinx)/(cosx) =1/cosx + tanx + 1/cosx - tanx =secx +secx =2secx

OpenStudy (mayaal):

got it....?

OpenStudy (anonymous):

THANK YOU!

OpenStudy (mayaal):

Aaisha where r u from...?

OpenStudy (anonymous):

what about this one? cot(x-pi/2)=-tanx any idea?

OpenStudy (anonymous):

from the US, wbu?

OpenStudy (mayaal):

I live on us but i was born in Pakistan.... :)

OpenStudy (mayaal):

in*

OpenStudy (mayaal):

let me see...

OpenStudy (mayaal):

cot(p/2-x)=tan x u knw that?

OpenStudy (anonymous):

no what?

OpenStudy (mayaal):

u see the cot function is an odd function so cot(-x)=-cot x

OpenStudy (mayaal):

that is the reversal happened here

OpenStudy (mayaal):

hence proved :)

OpenStudy (mayaal):

Got it...?